### v1

# MulSMo: Multimodal Stylized Motion Generation by Bidirectional Control Flow

Abstract
Generating motion sequences conforming to a target style while adhering to the given content prompts requires accommodating both the content and style. In existing methods, the information usually only flows from style to content, which may cause conflict between the style and content, harming the integration. Differently, in this work we build a bidirectional control flow between the style and the content, also adjusting the style towards the content, in which case the style-content collision is alleviated and the dynamics of the style is better preserved in the integration. Moreover, we extend the stylized motion generation from one modality, i.e. the style motion, to multiple modalities including texts and images through contrastive learning, leading to flexible style control on the motion generation. Extensive experiments demonstrate that our method significantly outperforms previous methods across different datasets, while also enabling multimodal signals control. The code of our method will be made publicly available.
1 Introduction
Stylized human motion generation is a rapidly growing field in computer vision, offering significant value across various applications like filmmaking, gaming, virtual reality, and robotics. The model is tasked to generate human motions according to instruction texts while conforming to a given style, which requires the model to accommodate both the content and style features.
Previous methods tackling this problem try to directly concatenate the two features of content and style [27], or integrate them by cross attention [21]. Recent methods borrow the ideas from stylized image generation, employing AdaIn [16], LoRA [15], or ControlNet [45] to control the style of the generated motion. AdaIn is an effective mechanism for enforcing global styles, yet it inherently has limitations when it comes to capturing local details of styles. LoRA has been successfully utilized to incorporate the target style into the generated content, but it requires training a separate LoRA model for each style, which is laborious for plentiful styles. In contrast, ControlNet exploits only one model to settle various detailed styles of the same mode, by taking in the control signal of the target style. Therefore, in this paper, we seek to push the limit of the ControlNet for stylized motion generation by designing a novel approach.
Figure 1: MulSMo enables multimodal signals to control the stylized motion generation.
Figure 2: Overview of MulSMo. Our approach can take in various modalities as style signals, such as style motion sequences, texts, and images. We generate stylized human motions by combining content text $\mathbf{c}$ with style signals $\mathbf{s}$. In the encoder part of the latent denoiser, we establish a bidirectional control flow between the generation network and the style network, utilizing respective zero linear for fusion. (Note that for simplicity we plot the style-to-content zero linear and content-to-style zero linear into one block in the figure.) This denoising step is repeated $T$ times to obtain the noise-free motion latent ${\bm{z}}_{0}$, which is decoded to stylized motion by the decoder.
While previous methods have achieved impressive results by adapting ControlNet for stylized motion generation [46], the control flow always moves from the style motion to the content motion. This brings a problem. The style motion often diverges from the content motion, leading to conflict in the integration of style and motion, such that the generation model tends to ignore the conflicted feature in the style. The conflict is especially severe in the motion dynamics, making it difficult to accurately transfer the dynamics of the style motion to the content motion. Therefore, the styles differing in dynamics would generate similar motions, such as ‘aeroplane’ and ‘flapping’. Another problem of previous methods lies in the control signals, which are limited to a style motion in previous methods [17, 13, 32, 46]. This leads to difficulties for users to input the target style, limiting the applications of stylized motion generation.
To address these two problems, in this paper, we propose a Multimodal Stylized Motion generation model (MulSMo), where a bidirectional control mechanism is designed to impose the style onto content motion, and the contrastive learning is leveraged to enable multi-modality style control, as shown in Figure 1. Specifically, we establish a bidirectional control flow between the style network and generation network, such that the control not only flows from the style to content, but also flows from the content to style, as displayed in Figure 2. This enables the style to also adjust towards the content, bridging the gap between two spaces and relieving the conflict, especially the divergence in dynamics. To enable both local and global bidirectional control, we equip the bidirectional flow modules in each encoder block so that local features can be controlled in the shallow layers and more global features is controlled while going deeper. To enable multimodal signals for style control, we propose to align the style embeddings across style motion sequences, text descriptions, and images within a contrastive learning manner. In this way, the style text descriptions or style images can play the same role as style motion sequences for style control during generation. Extensive experiments are performed with either style motion, texts, or images as style input, and demonstrate that our method obtains better performance compared to previous methods.
The main contributions of this work are then summarized as follows:
  • We novelly build a bidirectional motion ControlNet for stylized motion generation, where the control flow moves not only from the style to content but also from the content to style, in which case the style-content collision is relieved.
  • We extend the stylized motion generation from controlled by only style motion sequence to multimodal signals, where the style embeddings of different-modality signals are aligned by contrastive learning.
  • We outperform previous stylized motion generation and motion style transfer on multiple datasets, while enabling flexible style input from motions, texts, or images.
2 Related Work
2.1 Stylized Image Generation
Image style is generally represented by the global statistical characteristics of images. Early work by Gatys et al [10] demonstrates the feasibility of transferring visual styles through aligning Gram matrices within neural networks. Adaptive instance normalization (AdaIN) [16] is introduced to facilitate arbitrary image style transfer by dynamically adapting feature statistics. Recently, diffusion models [30] have achieved great success in image generation, and ControlNet [45] and LoRA [15] are widely work [19] used to control these models to generate stylized images. Specifically, ZipLoRA [31] allows for the generation of any subject in any style by effectively combining independently trained style and subject LoRAs. StyleAdapter [37] introduces a unified stylized image generation model capable of producing a variety of stylized images without the need for test-time fine-tuning.
2.2 Motion Generation Model
Human motion generation based on transformer models [22, 12] and diffusion models [8, 5, 2] has made significant progress in recent years. Momask [12] employs a motion residual VQ-VAE with multiple codebooks, enhancing generation through a residual transformer. LaMP [22] designs a novel motion-aware text encoder instead of CLIP to enhance the motion generation quality and proposes a motion-to-text large language model. MDM [35] establishes a solid baseline for text-to-motion generation and serves as a robust pretrained model for novel conditional motion generation. MLD [6] further enhances efficiency by embedding the diffusion process within the latent space. Inspired by the efficacy of diffusion models for control and conditioning, several works have controlled pretrained motion diffusion models to follow constraints such as trajectories [41, 20, 36], 3D scenes [43, 4], and object interactions [26, 39, 7].
2.3 Stylized Motion Generation
Motion style transfer [1, 25, 17, 34, 24, 32, 28, 21, 27, 33, 38, 42] is the most widely used technique to create stylized motion, transferring the style from a reference motion to a source motion. Aberman et al. [1] introduce a generative adversarial network to disentangle motion style from content and facilitate their re-composition, eliminating the need for paired data. Motion Puzzle [17] designs a generative framework that can control the style of individual body parts. Guo et al.[13] utilizes the latent space of pretrained motion models to improve the feature extraction and integration of motion content and style. Song et al.[32] proposes a diffusion-based approach that integrates trajectory awareness, and achieves style transfer through AdaIN.
However, a major limitation of the above models is that they rely on motion sequences to extract content information, while stylized motion generation [3, 46] can generate motions according to given texts while conforming to a target style. Inspired by the ControlNet, SMooDi [46] enables stylized motion generation from content text descriptions and style motion sequences by introducing two kinds of style guidance to steer the pretrained text-to-motion model toward the target style, realizing fine-grained style control. Nonetheless, the control flow only moves from the style to the content, leading to severe conflict in some cases. In contrast, we propose a bidirectional control mechanism to adapt the style towards content, alleviating the conflict.
3 Method
3.1 Overview
In this section, we present our proposed MulSMo, which generates stylized motion sequences by taking text descriptions and multi-modal style signals as input. An overview of our model is depicted in Figure 2. Aligning with the settings in MLD [6], we also place the diffusion process in the motion latent space. In our method, $\epsilon_{\theta}$ denotes the latent denoiser and $\{{\bm{z}}_{t}\}_{t=0}^{T}$ represents the noised latent sequences, where ${\bm{z}}_{T}$ is a Gaussian noise. Our purpose is to denoise from ${\bm{z}}_{T}$ to a clean stylized latent motion sequence. Given a content prompt $\mathbf{c}$ and a style signal $\mathbf{s}$, we define the denoising process at step $t\in(0,T]$ as $\epsilon_{t}=\epsilon_{\theta}({\bm{z}}_{t},t,\mathbf{c},\mathbf{s})$, where $\epsilon_{\theta}$ denotes the predicted noise at timestep $t$. After denoising for $T$ timesteps, we obtain a clean latent ${\bm{z}}_{0}$, which is then decoded by a pretrained decoder into a stylized motion sequence $\hat{{\bm{x}}}_{0}\in\mathbb{R}^{L\times D}$, where $L$ denotes the frame length of motion and $D$ is the dimension of human motion representations. Following [46], we adopt the same motion representations as detailed in HumanML3D [11], with $D=263$. As depicted in Figure 2, the text description $\mathbf{c}$ is used to control the content domain of the generated motion, while the style domain is controlled by style signals, including reference motion sequences, texts, or images.
Figure 3: Architectures with different control flow between generation network and style network. Based on experimental validation, we select (b) as the final architecture.
3.2 Bidirectional Motion Control Flow
We improve the ControlNet [45] by introducing a bidirectional feedback mechanism to make the transferred motion style more accurate. As depicted in Figure 3(a), existing works [46] directly inject the control signal into the generation network encoder. The conflict between the style and content features makes the generation unstable. In contrast, as shown in Figure 3(b), we build bidirectional fusion blocks between the style and content encoder to bridge the gap and enable more stable generation. The main idea is that the generation network and the control network engage in a high-frequency interaction, which synchronously adjusts the style and content features in a dynamic manner, minimizing the conflict and maintaining the dynamism of the style feature instead of being a simple pose.
As illustrated in the middle part of Figure 2, we build bidirectional fusion modules between the style and content encoders to bridge the information gap and resolve the conflict dynamically. To enable both local and global information to be shared between the two encoders, we design bidirectional fusion modules in each encoder block so that local information can be shared in the shallow layers and more global information is exchanged while going deeper. Denote $\mathcal{S}=\{\mathcal{S}_{i},i\in\{1,2,...,N\}\}$ the $N$ blocks style encoder, where $\mathcal{S}_{i}$ is the $i_{th}$ block and denote its output encoded feature as $F_{s_{i}}$. Similarly, we denote $\mathcal{G}=\{\mathcal{G}_{i},i\in\{1,2,...,N\}\}$ the content encoder, $\mathcal{G}_{i}$ the $i_{th}$ block and $F_{c_{i}}$ the encoded feature of the $i_{th}$ block. Denote fusion modules from style to contents as $\mathcal{L}_{s2c}=\{\mathcal{L}_{s2c_{i}},i\in\{1,2,...,N\}\}$. Then the fusion from the style to the content feature at the $i_{th}$ block can be denoted as:
$F_{s2c_{i}}=\mathcal{L}_{s2c_{i}}(F_{s_{i-1}})+F_{c_{i-1}}.$ (1)
Same as the style to content fusion, the content to style fusion modules $\mathcal{L}_{c2s}=\{\mathcal{L}_{c2s_{i}},i\in\{1,2,...,N\}\}$ performs fusion at the $i_{th}$ block as:
$F_{c2s_{i}}=\mathcal{L}_{c2s_{i}}(F_{c_{i-1}})+F_{s_{i-1}}.$ (2)
The fused features are then regarded as the output feature of the $i_{th}$ block as $F_{c_{i}}=F_{s2c_{i}}$ and $F_{s_{i}}=F_{c2s_{i}}$. They are then fed into the following blocks for further feature extraction and bidirectional fusion in a similar way. In the implementation, we use linear layers to perform fusion modules ($\mathcal{L}_{s2c}$ and $\mathcal{L}_{c2s}$) and initialize them with zero weights following LoRA [15] for stable training.
The designed bidirectional motion control flow bridges the information gap and synchronously resolves the conflict between the style and content. This is crucial for stable stylized motion generation, especially when the content and the style are in different domains, e.g., using texts or images for style control, which we will discuss in the following section.
Figure 4: Contrastive learning for enabling multi-modality stylized motion control. The adaptor can partially alleviate the gap between style motions and images/texts.
3.3 Multimodal Style Control
Previous methods, e.g., [3, 46] require style motion sequences for stylized motion control. However, capturing such style motion sequences is time-consuming and restricts these methods for scaling up. In this work, we enable multi-modality stylized motion control so that the generated motion can be easily stylized by text or image control, which is easier to obtain and is more user-friendly. To achieve this, our main idea is to align the feature between the style motion sequences, text descriptions, and the style images by contrastive learning. Specifically, we first utilize CLIP [29] pretrained from numerous image-text pairs to align the features between text descriptions and images. Then the aligned image-text features, after being projected by an adapter, are aligned with the style motion sequence feature within a contrastive learning scheme.
As shown in Figure 4, we input the style label into CLIP to obtain text embeddings $t$. After passing through an adaptor, we align the text embeddings $t$ with the style embeddings $s$ and optimize them using the InfoNCE loss [14]:
$\mathcal{L}_{\text{InfoNCE}}=-\log\frac{\exp(\text{sim}(t_{i},s_{j})/\tau_{t})}{\sum_{k=1}^{N}\exp(\text{sim}(t_{i},s_{k})/\tau_{t})},$ (3)
where $t_{i}$ and $s_{j}$ are positive samples, while $t_{i}$ and $s_{k}$ are negative samples. sim($\cdot$) means cosine similarity and $\tau$ denotes the temperature hyperparameter. During the inference stage, we can use a single label word as a textual control signal for stylized motion generation. Since CLIP has already aligned language and vision, we can also input images as control signals, which are processed through the adaptor to obtain style-aware image embeddings as style control.
In this way, we alleviate the domain gaps between image, text, and style motion sequence when serving as style control signals. These extracted control features are then fed into the bidirectional motion control flow modules for stylized motion generation. The proposed bidirectional control modules further resolve the conflicts between different domains and enable stable stylized motion generation given only text or image as style control signals.
Figure 5: Qualitative results of using text or image as the style control signal.
3.4 Stylized Motion Generation Guidance
Following [46], we adopt classifier-free guidance and classifier guidance to further enhance the stylized motion generation process. The combination of these two guidance methods makes the generated motion more consistent with the content prompt and the style signal.
Classifier-free Style Guidance. We divide the conditioned classifier-free guidance into two parts.
$\mathbf{\epsilon}_{\theta}({\bm{z}}_{t},t,\mathbf{c},\mathbf{s})=\,$ $\mathbf{\epsilon}_{\theta}({\bm{z}}_{t},t,\emptyset,\emptyset)$
$+$ $w_{c}\big{(}\mathbf{\epsilon}_{\theta}({\bm{z}}_{t},t,\mathbf{c},\emptyset)-\mathbf{\epsilon}_{\theta}({\bm{z}}_{t},t,\emptyset,\emptyset)\big{)}$
$+$ $w_{s}\big{(}\mathbf{\epsilon}_{\theta}({\bm{z}}_{t},t,\mathbf{c},\mathbf{s})-\mathbf{\epsilon}_{\theta}({\bm{z}}_{t},t,\mathbf{c},\emptyset)\big{)},$ (4)
where $w_{c}$ and $w_{s}$ represent the strengths of the classifier-free guidance for the condition $\mathbf{c}$ and $\mathbf{s}$, respectively. We slightly abuse the notations here by using $\emptyset$ to denote a condition is not used.
Classifier-based Style Guidance. To further improve the stylization of a motion diffusion model, we adopt the classifier guidance [9, 44] to provide stronger guidance to the generated motion towards the desired style. The core of our classifier-based style guidance is a novel analytic function $G({\bm{z}}_{t},t,\mathbf{s})$, which calculates the $L_{1}$ distance between the style embedding of the generated clean motion $\hat{{\bm{x}}}_{0}$ at denoising step $t$ and the reference style motions $\mathbf{s}$. The function’s gradient is utilized to steer the generated motion towards the target style.
$\mathbf{\epsilon}_{\theta}({\bm{z}}_{t},t,\mathbf{c},\mathbf{s})$ $=\mathbf{\epsilon}_{\theta}({\bm{z}}_{t},t,\mathbf{c},\mathbf{s})+\tau\nabla_{{\bm{z}}_{t}}G({\bm{z}}_{t},t,\mathbf{s}),$
$G({\bm{z}}_{t},t,\mathbf{s})$ $=|f(\hat{{\bm{x}}}_{0})-f(\mathbf{s})|,$ (5)
where $\tau$ adjusts the strength of reference-style guidance and $f$ denotes the style feature extractor. The generated motion $\hat{{\bm{x}}}_{0}$ is obtained by first converting the denoising output latent ${\bm{z}}_{t}$ into the predicted clean latent as shown below:
$\hat{{\bm{z}}}_{0}=\frac{{\bm{z}}_{t}-\sqrt{1-\alpha_{t}}\varepsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}},{\bm{s}})}{\sqrt{\alpha_{t}}},$ (6)
where $\alpha_{t}$ denotes the pre-defined noise scale in the forward process of the diffusion model. The predicted clean latent $\hat{{\bm{z}}}_{0}$ is then input into the motion decoder $\mathbf{D}$ to obtain the generated motion.
The style feature extractor is obtained by training a style classifier on the 100STYLE dataset [24] and removing its final layer. Training the style feature extractor with ground-truth style labels for supervision enables it to effectively capture style-related features. Consequently, style classifier guidance can provide more precise support for stylized motion generation.
Figure 6: Qualitative results of stylized motion generation on the 100STYLE dataset. The style input is a style motion while the content input is a text prompt. Note that the ‘Flapping’ style involves the up-and-down motion.
[…]
(Top 3) Ratio $\downarrow$
Motion MLD+Aberman [1] 54.367 3.309 5.983 0.406 8.816 0.347
MLD+Motion Puzzle [17] 63.769 6.127 6.467 0.290 6.476 0.185
SMooDi [46] 72.418 1.609 4.477 0.571 9.235 0.124
Ours 77.042 1.527 4.292 0.613 9.303 0.118
SMooDi [46] 81.298 3.066 5.011 0.439 9.433 0.155
Ours 86.455 3.028 4.872 0.498 9.514 0.153
Text ChatGPT+MLD [46] 4.819 0.614 4.313 0.605 8.836 0.131
Adaptor+MLD (Ours) 52.100 0.754 3.875 0.674 8.875 0.104
Table 1: Quantitative comparison with baseline methods of stylized motion generation on 100STYLE dataset. We utilize style motion and text as control signals respectively. Note that there is a tradeoff between style reflection metrics (SRA) and content preservation metrics (R-precision, FID, MM Dist). SMooDi and Ours indicate that we adjust $w_{s}$ to improve the style intensity and reach higher SRA. The best results under the same setting are in bold.
3.5 Learning Scheme
We freeze the parameters of MLD and only train the style network on the 100STYLE dataset using the following loss function:
$\mathcal{L}_{std}=\mathbb{E}_{\epsilon,{\bm{z}}}\left[\left\|\epsilon_{\theta}({\bm{z}}_{t},t,\mathbf{c},\mathbf{s})-\epsilon\right\|^{2}_{2}\right],$ (7)
where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ represents the ground-truth noise added to ${\bm{z}}_{0}$. Experimentally, when training solely on the 100STYLE dataset, we observe a phenomenon similar to that mentioned in SMooDi [46]: the model gradually tends to resample the reference motions rather than following the given content text. Therefore, we adopt a content prior preservation loss $L_{pr}$, on the HumanML3D dataset:
$\mathcal{L}_{pr}=\mathbb{E}_{\epsilon^{\prime},{\bm{z}}^{\prime}}\left[\left\|\epsilon_{\theta}({\bm{z}}_{t}^{\prime},t,\mathbf{c}^{\prime},\mathbf{s}^{\prime})-\epsilon^{\prime}\right\|^{2}_{2}\right],$ (8)
where ${\bm{z}}_{t}^{\prime}$, $\mathbf{c}^{\prime}$ and $\mathbf{s}^{\prime}$ represents the motion latent, content prompt and style motion sequence derived from the HumanML3D dataset. $\epsilon^{\prime}$ is the noise map added to ${\bm{z}}^{\prime}_{0}$. With content prior preservation loss, our model can maintain the ability to follow the content text constraint while learning diverse motion styles from the 100STYLE dataset.
Overall, the training loss function of our framework is defined as follows:
$\mathcal{L}_{all}=\mathcal{L}_{std}+\lambda_{pr}\mathcal{L}_{pr},$ (9)
where $\lambda_{pr}$ and $\lambda_{cyc}$ are hyperparameters. We refer readers to the pseudocode and illustration for training in the supplementary material for more details.
4 Experiments
4.1 Experiment Setup
Implementation Details.
We utilize MLD as the pretrained generation network and train the style network from scratch. The style motion encoder in Figure 4 consists of a single transformer encoder, and the adaptor is a simple MLP. During the training of the style control network, the style network composed of four Transformer Encoder blocks is optimized while the parameters of the MLD are frozen. The framework is trained with the AdamW optimizer [23] and the learning rate is 1e-5. More details can be found in the supplementary.
Datasets.
We utilize the HumanML3D dataset [11] as the motion content dataset and the 100STYLE dataset [24] as the motion style dataset for training. Moreover, we use Xia dataset [40] as the style dataset for evaluating the motion style transfer task. The HumanML3D dataset is the largest motion content dataset paired with text annotations and is widely used for text-to-motion generation. The 100STYLE dataset [24] is the largest motion style dataset. Xia dataset is a smaller motion style collection encompassing 8 distinct styles with precise annotations for 8 action types. The motions within this collection are generally under 3 seconds in duration. Due to differences in skeletons between the 100STYLE dataset and HumanML3D, we retarget the motions from 100STYLE to match the HumanML3D (SMPL) skeleton. Following this alignment, we apply the same processing steps as used for the HumanML3D dataset to preprocess the 100STYLE dataset. We also follow SMooDi [46] and use MotionGPT [18] to generate pseudo text descriptions for the motion sequences. For both the HumanML3D dataset and the 100STYLE dataset, we use the consistent root-velocity motion representations outlined in [11].
Evaluation Metrics.
The metrics are designed to assess three key dimensions: Content Preservation, Style Reflection, and Realism. For evaluating content preservation, we follow [6] and use metrics including motion-retrieval precision (R-precision), Multi-modal Distance (MM Dist), Diversity, and Frechet Inception Distance (FID). For realism, we address common foot skating issues by integrating the Foot Skating Ratio metric proposed by [20]. For style reflection, we follow SMooDi [46] to use Style Recognition Accuracy (SRA) [17]. SRA and FID are two metrics that involve a trade-off. A higher SRA indicates more pronounced stylization but leads to a relatively worse FID. During the evaluation, a content text from the HumanML3D dataset and a motion style sequence from the 100STYLE dataset, are randomly selected to generate the stylized motion. A pretrained style classifier then computes the SRA for the generated motion, following previous methods [46].
[…]
Ratio $\downarrow$
MLD+Aberman [1] 61.006 3.892 0.338
MLD+Motion Puzzle [17] 67.233 6.871 0.197
SMooDi [46] 65.147 1.582 0.095
Ours 70.238 1.566 0.087
Table 2: Quantitative results of motion style transfer on the HumanML3D dataset.
[…]
Ratio $\downarrow$
MLD+Aberman [1] 56.667 34.444 5.681 0.0260
MLD+Motion Puzzle [17] 67.778 25.556 5.360 0.0316
SMooDi [46] 61.111 45.555 4.663 0.0317
Ours 67.870 47.071 4.582 0.0304
Table 3: Quantitative results of motion style transfer on the Xia dataset. We report the Content Recognition Accuracy (CRA).
[…]
(Top 3) ratio $\downarrow$
(a) 72.418 1.609 4.477 0.571 9.235 0.124
(b) (Ours) 77.042 1.527 4.292 0.613 9.303 0.118
(c) 78.623 3.620 5.766 0.393 7.872 0.142
(d) 51.527 4.869 5.118 0.508 6.747 0.129
Table 4: Ablation study on different control flow. For (b), (c), and (d), we utilize the same guidance scale and style guidance scale. In contrast, (a) employs the optimal guidance scale and style guidance scale as described in the SMooDi.
Baselines.
We focus on three primary tasks: motion-guided stylized motion generation, motion style transfer, and text/image-guided stylized motion generation. For the motion-guided stylized motion generation and motion style transfer tasks, we utilize the evaluation metrics established in SMooDi [46], incorporating MLD as the text-to-motion model; the methods for motion style transfer are consistent with those employed in the motion style transfer task. For the text/image-guided stylized motion generation task, we extract features from the texts and images using CLIP and input these features into our trained adaptor as style embeddings for style control. Unlike ChatGPT+MLD, which essentially functions as a straightforward text-to-motion model without embodying the control aspect, our approach allows for style control using a simple label word, enabling more nuanced manipulation of the generated motion styles.
4.2 Evaluation
Motion-guided stylized motion generation.
For the stylized text-to-motion generation task with motions as style, we present a comparative analysis of our approach against four baseline methods, as summarized in Table 1. When we input style motion as the control signal, our model outperforms existing methods on both SRA and FID metrics. SRA and FID are two trade-off metrics; however, compared to SMooDi [46], our approach achieves an FID that is lower by 5.1% and 1.3% while maintaining an SRA that is 6.4% higher. As illustrated in Figure 6, SMooDi exhibits conflicts between style and content, resulting in the stylized motion failing to align with the content prompts. Furthermore, the style motion tends to lose some dynamic qualities, leading to inaccuracies in classifying stylized actions for similar poses, such as ‘Flapping’ and ‘Aeroplane’.
Text/image-guided stylized motion generation.
When the control signal comes to texts, ChatGPT+MLD utilizes ChatGPT to merge style labels from 100STYLE with texts from HumanML3D into a sentence. For example, ‘a person walks’ + ‘old’ = ‘an elderly person walks’. Then this merged sentence is fed to MLD. The performance of ChatGPT+MLD is limited, achieving merely 4.819% in the SRA metric. This suggests that the MLD framework, despite incorporating style descriptors, falls short in facilitating stylized generation from text inputs alone. In contrast, our Adaptor+MLD accepts one label word as signal control, and fully leverages the control mechanism of the MulSMo, improving the SRA performance from 4.819% to 52.1% with only a small degradation on the FID. The visualization of the generated stylized motions from texts or images is displayed in Figure 5, from where we can see the convenient multimodal control.
Motion style transfer.
For the motion style transfer task, we follow the settings of SMooDi and evaluate only three metrics: FID, SRA, and Foot skating ratio. We utilize the HumanML3D dataset [11] as the source of motion content and extract motion styles from the 100STYLE dataset. As shown in Table 2, our algorithm achieves a significant improvement of 7.8% on the SRA metric, along with advancements of 1.0% and 8.4% on the FID and Foot skating ratio metrics, respectively. Furthermore, we conduct experiments on the Xia dataset [40], which was not seen during the training phase of both our model and the baseline model. Given the presence of motion content labels in the Xia dataset, we report the Content Recognition Accuracy (CRA). The results are presented in Table 3.
Figure 7: User study on the stylized motion generation task.
4.3 Ablation Study
To validate the effectiveness of our framework, we conduct ablation studies on different control-feedback flows. As shown in Figure 3, we evaluate four architectures: (a), (b), (c), and (d). The ablation results are shown in Table 3.
In Figure 3 (a), we align our approach with SMooDi, introducing a unidirectional control signal between the encoder of the generation network and the encoder of the style network, with results corresponding to the first row in Table 4. In Figure 3 (b), we implement a bidirectional control flow between these two encoders, where the style signal influences the content, and at the same time, the content signal provides feedback to the style. We choose this as the final framework. In Figures 3 (c) and 3 (d), bidirectional signals are also incorporated between the middle block and the decoder, respectively. These two frameworks lead to a degradation in the final results. We attribute this to the control signals from the style network disrupting the parameters of the frozen middle block and decoder of the generation network, thereby affecting the recovery of fine details.
4.4 User Study
In order to address the inherently subjective nature of stylized motion, we conduct user studies employing pairwise comparisons to systematically evaluate our method in the task of stylized motion generation. A total of 41 human subjects were recruited to participate in this study. During each trial, each participant will be provided with two video segments, each approximately 9 seconds in length, generated by our method and SMooDi, respectively. They are tasked with selecting their preferred video, taking into account three key dimensions: Realism, Style Reflection, and Content Preservation. As illustrated in Figure 7, our method garnered significantly higher user preference across all three dimensions when compared to the SMooDi. Details regarding the study are provided in the supplementary materials.
5 Conclusion
This paper proposes a multimodal stylized motion generation framework that innovatively leverages style motion, text, or image as style control to produce dynamic motion sequences. Different from previous methods where the control information only moves from style to content, a bidirectional control flow between the style network and the generation network is also introduced to resolve the conflict between the two spaces for better stylized motion generation.
Limitations
Although the conflict between the content and style is relieved, the level of similarity between the generated motion and the style, and that between the generated motion and the content text is still a trade-off.